# Optimization Insights into Deep Diagonal Linear Networks

### v2

@@ Abstract
Overparameterized models trained with (stochastic) gradient descent are ubiquitous in modern machine learning. These large models achieve unprecedented performance on test data, but their theoretical understanding is still limited. In this paper, we take a step towards filling this gap by adopting an optimization perspective. More precisely, we study the implicit regularization properties of the gradient flow “algorithm” for estimating the parameters of a deep diagonal neural network. Our main contribution is showing that this gradient flow induces a mirror flow dynamic on the model, meaning that it is biased towards a specific solution of the problem depending on the initialization of the network. Along the way, we prove several properties of the trajectory.
@@ 1 Introduction
In recent years, the application of deep networks has revolutionized the field of machine learning, particularly in tasks involving complex data such as images and natural language. These models, typically trained using stochastic gradient descent, have demonstrated remarkable performance on various benchmarks, raising questions about the underlying mechanisms that contribute to their success. Despite their practical efficacy, the theoretical understanding of these models remains relatively limited, creating a pressing need for deeper insights into their generalization abilities. The classical theory shows that the latter is a consequence of regularization, which is the way to impose a priori knowledge into the model and to favour “simple” solutions. While usually regularization is achieved either by choosing simple models or explicitly adding a penalty term to the empirical risk during training, this is not the case for deep neural networks, which are trained simply by minimizing the empirical risk. A new perspective has then emerged in the recent literature, which relates regularization directly to the optimization procedure (gradient based methods). The main idea is to show that the training dynamics themselves exhibit self regularizing properties, by inducing an implicit regularization (bias) which prefers generalizing solutions (see [Vardi, 2023] for an extended review of the importance of implicit bias in machine learning).
In this context, a common approach is to study simplified models that approximate the networks used in practice. Analyzing the implicit bias of optimization algorithms for such networks is facilitated but still might give some insights on the good performance of neural networks in various scenarios.
1.1 Related works
Matrix factorization
A step in this direction is made by [Gunasekar et al., 2017] who analyze the implicit regularization induced by matrix factorization for minimizing a simple least squares objective as $\mathcal{L}\left(\Theta\right)=\|A\Theta-b\|^{2}$ for $\Theta\in\mathbb{R}^{n\times n}$. More precisely, the authors consider the reparametrization $\Theta=UU^{T}$ for $U\in\mathbb{R}^{n\times d}$ and so the new problem
$\min_{U\in\mathbb{R}^{n\times d}}\left\|A\left(UU^{T}\right)-b\right\|^{2}.$ (1)
Under some assumptions on the initialization, they demonstrate that using gradient flow on the variable $U$ returns a matrix $\Theta=UU^{T}$ approximating a specific solution of the original problem, the one with minimal nuclear norm. In other words, a simple change of variable before running the algorithm allows to promote sparsity of the recovered solution, which tends to imply a better generalization. Such a setting can roughly be linked to neural networks and has been followed by several works enriching the understanding of training factorized matrices: [Li et al., 2018] study the optimization aspects of such a model; [Arora et al., 2018] show that such a reparameterization can improve the conditioning of the problem, accelerating the resulting optimization problem; [Gunasekar et al., 2018a] give implicit bias results for an overparameterized matrix factorization; [Arora et al., 2019] study the effect of deep matrix factorization, proving a similar implicit bias towards a minimal nuclear norm solution for sufficiently small initialization.
Weight normalization
Similarly, weight normalization is a strategy used in machine learning that has been studied using this framework. It has been proven to accelerate training ([Salimans and Kingma, 2016]) and to have desirable implicit regularization properties ([Wu et al., 2020]). Moreover, it can be calibrated to take full advantage of both characteristic ([Chou et al., 2024]).
More complex parameterizations
Several works focus on more complex architectures, showing valuable but less interpretable features; for instance, studying the implicit bias in the case of linear convolutional networks ([Gunasekar et al., 2018b]), shallow neural networks ([Allen-Zhu et al., 2019a, Allen-Zhu et al., 2019b]), ReLU networks ([Vardi and Shamir, 2021]) and deep linear networks ([Marion and Chizat, 2024]). We can also mention the work of [Chizat and Bach, 2020], where a two-layer neural network with infinite width is considered.
Diagonal Linear Networks
Diagonal Linear Networks are introduced in the literature by [Woodworth et al., 2020] and [Moroshko et al., 2020] as a linear approximation of a neural network with diagonal connections between the nodes (see Figure 1). In this case the network function is given by $f_{\theta}(x)=\theta^{T}x$ with the parameters vector expressed as $\theta=u\odot v$, where $\odot$ denotes the Hadamard product. Although this model appears to be overly simplistic, its analysis offers a considerable degree of insight. Both [Woodworth et al., 2020] and [Moroshko et al., 2020] highlight the role of the initialization for such a network in regard of the implicit bias induced on the solution: a small initialization favours a solution close to that with minimal $L^{1}$ norm, while a large one tends to minimize the $L^{2}$ norm. These two regimes can be put in parallel with respectively the rich regime and the kernel (or lazy) regime outlined in [Chizat et al., 2019]. [Azulay et al., 2021] give additional results on the implicit bias of the network, allowing untied weights and also considering the case of connected networks (non diagonal). This model has been also been analyzed to better understand the implicit bias induced by stochastic gradient descent [Pesme et al., 2021, Even et al., 2023], by the choice of the step size [Nacson et al., 2022] and by other methods involving momentum [Papazov et al., 2024]. The $L$-layer case has been studied by [Woodworth et al., 2020, Moroshko et al., 2020, Chou et al., 2023a] for a very simplified model, suggesting that the number of layers is positively correlated with the sparsity of the selected solution.
@@ 1.1 Related works
Diagonal Linear Networks are closely related to Hadamard parameterization (HP) and it is worth noticing that HP was used before for sparsity recovery. [Hoff, 2017] takes benefit of HP to promote sparsity in a LASSO problem. [Vaskevicius et al., 2019, Zhao et al., 2022] also consider an HP for a least squares problem inducing an implicit bias towards the minimal $L^{1}$-norm solution, enabling the use of early stopping strategies. In a comparable vein, [Amid and Warmuth, 2020] exploit a reparameterization $\theta=u^{\odot 2}$ for a classification problem and [Chou et al., 2022] apply HP to solve non negative least squares. [Poon and Peyré, 2023] elegantly reformulate a group-LASSO problem using some Hadamard parameterization in order to apply more efficient optimization methods. The geometry induced by HP is also a subject of interest and in this regard, [Ouyang et al., 2024] give necessary condition for a reparameterized function to satisfy the Kurdyka-Łojasiewicz property.
@@
As shown in [Azulay et al., 2021, Chou et al., 2023a], Diagonal Linear Networks trained with a Gradient Flow (or Gradient Descent) share direct links with Mirror Flow (or Mirror Descent). One of the first appearances of the Mirror Flow dynamic can be attributed to [Alvarez et al., 2004]. Many years later, a number of studies are presented in the context of understanding the effects of reparameterization. The mirror gradient method having an implicit bias towards specific solutions, [Vaskevicius et al., 2020] propose early stopping strategies taking advantage of it. [Li et al., 2022] provide assumptions that are necessary for a reparameterization trained with a Gradient Flow to be equivalent to a Mirror Flow. In their paper, [Chou et al., 2023b] adopt an original point of view, seeing reparameterization as a way to enforce some implicit bias. In this perspective, the authors give guidelines to define it efficiently.
1.2 Contributions
The general problem stated at the beginning of the introduction can be phrased in a precise mathematical form. We want to study the properties of an optimization procedure run on the following problem, that is the problem available in practice (e.g. an empirical risk depending on some parameters):
$\min_{u\in\mathbb{R}^{p}}\ \tilde{\mathcal{L}}(u).$ (2)
We look at this through the lens of reparameterization: we suppose that the previous problem arises from a hidden problem through some change of variable, i.e. that there exists a loss function $\mathcal{L}:\mathbb{R}^{d}\rightarrow\mathbb{R}$ and a reparametrization $q:\mathbb{R}^{p}\to\mathbb{R}^{d}$ such that
~$\tilde{\mathcal{L}}(u)=\mathcal{L}(q(u)).$
In the variable $\theta=q(u)$, the hidden problem is then given by
$\min_{\theta\in\mathbb{R}^{d}}\ \mathcal{L}(\theta).$ (3)
The general question of this line of research is the following. Running an optimization procedure on the reparametrization $u\in\mathbb{R}^{p}$ to minimize the loss $\tilde{\mathcal{L}}$ of the available problem in 2, what happens on the variable $\theta=q(u)$ in terms of the minimization of $\mathcal{L}$, the hidden problem in 3?
One example of the previous setting is given by Shallow Linear (connected) Neural Networks, where, for $u=(W,v)$, $q(W,v)=W^{T}v$.
In this paper we focus on a Deep Diagonal Linear Networks, an extension of Diagonal Linear Networks to the $L$-layer case. The corresponding reparameterization of $\theta\in\mathbb{R}^{d}$ is given by
$\theta=q(u^{1},u^{2},\dots,u^{L})=u^{1}\odot u^{2}\odot\dots\odot u^{L},$
where $L$ is the number of layers and $u^{i}\in\mathbb{R}^{d}$ for every $i=1,...,L$. In this case, considering for instance the empirical risk with linear-quadratic loss on input-output data $(x_{i},y_{i})_{i=1}^{n}$, we have
$\min_{(u^{1},u^{2},\dots,u^{L})}\quad\tilde{\mathcal{L}}(u^{1},u^{2},\dots,u^{L})=\mathcal{L}(q(u^{1},u^{2},\dots,u^{L}))=\sum_{i=1}^{n}(\langle u^{1}\odot u^{2}\odot\dots\odot u^{L}),x_{i}\rangle-y_{i})^{2},$
where $\mathcal{L}(\theta)=\sum_{i=1}^{n}(\langle\theta,x_{i}\rangle-y_{i})^{2}$.
Although simplified versions of this model have been previously studied in the literature (see [Woodworth et al., 2020, Moroshko et al., 2020, Chou et al., 2023a]), this formulation appears to better approximate real-life networks. In fact, these works consider the reparameterization $\theta=u^{\odot L}-v^{\odot L}$, allowing for the derivation of numerous properties. Nevertheless, it can be argued that this model is overly simplistic, as the number of layers does not affect the number of hyperparameters.
The objective of this work is to gain a deeper comprehension of the effect of reparameterization when the network layers $(u^{j})_{j\in[L]}$ are trained using Gradient Flow. In this regard, we make the following contributions:
  • We identify quantities that are conserved along the trajectory of the flow. This allows us to show that the dynamic of the variable $\theta$ is driven by that of an abstract layer containing the minimal nodes.
  • We provide a mild assumption on the initialization such that applying Gradient Flow to the layers $(u^{j})_{j\in[L]}$ generates a trajectory in the variable $\theta$ that is the solution of a Mirror Flow.
  • We show several convergence properties in terms of the original optimization problem, under the same assumption. In particular, we show that the value of the loss function along the variable $\theta$ decreases exponentially to its minimal value under a Polyak-Łojasiewicz property. These results also reveal that initializing the layers with small values leads to a slow training, a phenomenon that is known in the literature for other models.
@@ 1.3 Notation
For any positive integer $L$, the set of integers from $1$ to $L$ is denoted by $[L]$. The coordinate-wise product, also called the Hadamard product, of two vectors $x$ and $y$ in $\mathbb{R}^{p}$ is denoted $x\odot y$, where for each $i$ in $[p]$, $(x\odot y)_{i}=x_{i}y_{i}$. The vector $x^{\odot L}$ is defined as $\left(x^{L}_{i}\right)_{i\in[p]}$ and if $\left(x^{j}\right)_{j\in[L]}\in\left(\mathbb{R}^{p}\right)^{L}$, then $\bigodot_{j=1}^{L}x^{j}$ refers to $\left(\prod_{j=1}^{L}x^{j}_{i}\right)_{i\in[p]}$. We use ||$\cdot$|| ⋅ | to denote the absolute value which can be defined for vectors in $\mathbb{R}^{p}$: for any $x\in\mathbb{R}^{p}$, $|x|=\left(|x_{i}|\right)_{i\in[p]}$. Similarly, the sign function denoted $\text{sign}(\cdot)$, the square root function $\sqrt{\cdot}$, the hyperbolic sine $\sinh$, the inverse hyperbolic sine arcsinh and the logarithm $\log$ are considered as functions from $\mathbb{R}^{p}$ to $\mathbb{R}^{p}$ that apply to each component. The Jacobian of $G=\left(G_{1},\ldots,G_{p}\right)$ with respect to $\theta\in\mathbb{R}^{k}$ is denoted by $J_{G}(\theta)$ and is defined as:
@@
[…]
@@ 2 Deep Diagonal Linear Networks
In this section, we state several results on the trajectory of the variable $\theta$ when parameterized by a Deep Diagonal Linear Network and when its parameters are trained by Gradient Flow. More specifically, we overparameterize our system by defining
$\theta=\bigodot_{j=1}^{L}u^{j},$
where $L\geqslant 2$ is the number of layers of the network and each $u^{j}\in\mathbb{R}^{d}$ corresponds to the $j$-th layer, as illustrated in Figure 2. Note that for $L=2$, we recover the classical Diagonal Linear Network. We consider the training of the Deep Diagonal Linear Network by Gradient Flow, which can be written as follows:
$\forall j\in[L],\leavevmode\nobreak\ \forall t\geqslant 0,\quad\frac{du^{j}(t)}{dt}+\nabla_{j}\mathcal{L}\left(\bigodot_{k=1}^{L}u^{k}(t)\right)=0,$ (4)
where we denoted $\left(\nabla_{j}\mathcal{L}\left(\bigodot_{k=1}^{L}u^{k}\right)\right)_{i}=\frac{\partial}{\partial u^{j}_{i}}\mathcal{L}\left(\bigodot_{k=1}^{L}u^{k}\right)$ for any $j\in[L]$ and $i\in[d]$. By simple computations, we get the following dynamical system:
$\forall j\in[L],\leavevmode\nobreak\ \forall t\geqslant 0,\quad\frac{du^{j}(t)}{dt}+\left(\bigodot_{k\neq j}u^{k}(t)\right)\odot\nabla\mathcal{L}(\theta(t))=0.$ (5)
@@
We first derive several properties of the trajectories of the weights of the network. We then demonstrate the main result on the dynamic of $\theta$ by adopting a different approach. Convergence guarantees are stated in the last section, under the same assumption on the initialization.

### v3

@@ Abstract
Gradient-based methods successfully train highly overparameterized models in practice, even though the associated optimization problems are markedly nonconvex. Understanding the mechanisms that make such methods effective has become a central problem in modern optimization. To investigate this question in a tractable setting, we study Deep Diagonal Linear Networks. These are multilayer architectures with a reparameterization that preserves convexity in the effective parameter, while inducing a nontrivial geometry in the optimization landscape. Under mild initialization conditions, we show that gradient flow on the layer parameters induces a mirror-flow dynamic in the effective parameter space. This structural insight yields explicit convergence guarantees, including exponential decay of the loss under a Polyak–Łojasiewicz condition, and clarifies how the parametrization and initialization scale govern the training speed. Overall, our results demonstrate that deep diagonal over parameterizations, despite their apparent complexity, can endow standard gradient methods with well-behaved and interpretable optimization dynamics.
@@ 1 Introduction
In recent years, the application of deep neural networks has transformed the field of machine learning, achieving outstanding results in tasks involving complex data such as images and natural language. These models are typically trained successfully using gradient-based methods, most notably stochastic gradient descent, consistently reaching zero empirical loss across various benchmarks. This empirical success is particularly striking given the highly nonconvex nature of the underlying optimization problems, where classical theory would provide little/no guarantees. Understanding why such simple optimization procedures are so effective in training overparameterized models has therefore become a central research question.
A common approach to understanding optimization in overparameterized models is to analyze simplified settings where the interaction between parameterization and optimization becomes more transparent. Even in these reduced models, the dynamics often remain surprisingly intricate, yet they can shed light on how parameterization shapes convergence behavior. Within this line of work, Diagonal Linear Networks (DLNs), introduced independently by [37] and [26], have emerged as a particularly tractable framework. In shallow DLNs, the network function takes the form $f_{\theta}(x)=\theta^{\top}x$ with parameters $\theta=u\odot v$, where $\odot$ denotes the Hadamard product (see Figure 1). Despite their simplicity, DLNs display a variety of nontrivial optimization phenomena. Prior studies have examined how initialization scale, step size, or stochasticity influence convergence and the nature of the obtained solution [7, 30, 15, 27, 29]. These works highlight, among other insights, that mild overparameterization can already smooth the loss landscape, and that small versus large initializations give rise to qualitatively different training behaviors, mirroring the so called rich and lazy regimes of [9].
@@
In this paper, we extend this line of inquiry to Deep Diagonal Linear Networks, which generalize DLNs to an $L$-layer parametrization. In this model, the effective parameter is given by
$\theta=u^{1}\odot u^{2}\odot\dots\odot u^{L},$
where each $u^{j}\in\mathbb{R}^{d}$. This formulation offers a closer abstraction of multilayer architectures while preserving analytical tractability, and it allows us to explore how depth influences the optimization landscape and the implicit dynamics induced by training. Note that a simpler version of the $L$-layers case has been studied in [37, 26, 10] considering the simpler parameterization $\theta=u^{\odot L}-v^{\odot L}$.
Our goal is to understand the effect of this reparameterization when the layers weights $(u^{j})_{j\in[L]}$ are trained using gradient flow. More precisely, we consider the problem
$\min_{\theta\in\mathbb{R}^{d}}\mathcal{L}(\theta),$
for some differentiable loss function $\mathcal{L}:\mathbb{R}^{d}\rightarrow\mathbb{R}$, and for some input-output data $X:\mathbb{R}^{d}\rightarrow\mathbb{R}^{n}$ and $y\in\mathbb{R}^{n}$. We optimize $\mathcal{L}$ by applying gradient flow with respect to the variables $(u^{j})_{j\in[L]}$.
Our main contributions are summarized as follows:
  • We provide a mild initialization condition under which gradient flow on $(u^{j})_{j\in[L]}$ induces a trajectory in $\theta$ that follows a mirror-flow dynamic.
  • Under the same condition, we establish several convergence guarantees for the original optimization problem. In particular, when the loss function $\mathcal{L}$ satisfies a Polyak-Łojasiewicz condition, the objective function in the $\theta$-space decreases exponentially toward its minimum. Our analysis also indicates that very small initializations lead to slow training, consistent with phenomena observed in related models.
  • We identify quantities that are conserved along the gradient-flow trajectory, in line with recent works on conservative laws [22, 24]. In particular, this reveals that the dynamics of $\theta$ are governed by those of an abstract layer associated with minimal components.
We provide a detailed review of related topics in the literature in the next section.
1.1 Related works
Conservation laws
A common approach to studying the behavior of optimization dynamics in neural network models is to identify the conserved quantities that remain invariant during training. Such quantities often provide insight into convergence properties or implicit bias phenomena (see, e.g., [33, 14]). Characterizing these conservation laws is therefore crucial for understanding the overall training dynamics. Recent works have pursued this direction by establishing convergence laws for gradient flow in general neural network models [22], momentum-based dynamics [23], and discrete-time dynamics [24].
@@ 1.1 Related works
Diagonal Linear Networks are closely related to Hadamard parameterization (HP) and it is worth noticing that HP was used before for sparsity recovery. [19] takes benefit of HP to promote sparsity in a LASSO problem. [35, 40] also consider an HP for a least squares problem inducing an implicit bias towards the minimal $L^{1}$-norm solution, enabling the use of early stopping strategies. In a comparable vein, [4] exploit a reparameterization $\theta=u^{\odot 2}$ for a classification problem and [12] apply HP to solve non negative least squares. [31] elegantly reformulate a group-LASSO problem using some Hadamard parameterization in order to apply more efficient optimization methods. The geometry induced by HP is also a subject of interest and in this regard, [28] give necessary condition for a reparameterized function to satisfy the Kurdyka-Łojasiewicz property.
@@
As shown in [7, 10], Diagonal Linear Networks trained with a Gradient Flow (or Gradient Descent) share direct links with Mirror Flow (or Mirror Descent). One of the first appearances of the Mirror Flow dynamic can be attributed to [3]. Many years later, a number of studies are presented in the context of understanding the effects of reparameterization. The mirror gradient method having an implicit bias towards specific solutions, [36] propose early stopping strategies taking advantage of it. [21] provide assumptions that are necessary for a reparameterization trained with a Gradient Flow to be equivalent to a Mirror Flow. In their paper, [11] adopt an original point of view, seeing reparameterization as a way to enforce some implicit bias related to the corresponding mirror map. In this perspective, the authors give guidelines to define it efficiently.
Matrix factorization
A significant step toward understanding the optimization dynamics of overparameterized models is provided by [18], who analyze gradient-based training in a simple least squares setting,
$\mathcal{L}(\Theta)=\|A\Theta-b\|^{2},\quad\Theta\in\mathbb{R}^{n\times n}.$
By introducing the factorization $\Theta=UU^{\top}$ with $U\in\mathbb{R}^{n\times d}$, the problem becomes
$\min_{U\in\mathbb{R}^{n\times d}}\|A(UU^{\top})-b\|^{2}.$
Under mild assumptions on initialization, the authors show that gradient flow on $U$ converges to a factorization corresponding to the minimal nuclear norm solution of the original problem. This result illustrates how reparameterization fundamentally changes the optimization landscape, leading gradient descent to converge toward more ”regularized” solutions. Subsequent works have deepened this understanding: [20] analyze the algorithmic regularization effects arising in such factorized settings; [5] demonstrate that overparameterization can improve conditioning, thereby accelerating convergence; [17] explore the geometry of the optimization path itself; and [6] extend the analysis to deep matrix factorization, linking initialization scale to convergence toward minimal nuclear norm solutions.
Weight normalization
Weight normalization, first introduced by [32], can also be viewed through the lens of optimization dynamics. By decoupling the weight direction and scale, this reparameterization improves the conditioning of the optimization problem and often accelerates convergence. Later analyses, such as [38], formalize its convergence properties and describe how the induced geometry affects the optimization trajectory. More recently, [13] have shown that weight normalization can be calibrated to simultaneously improve robustness and optimization efficiency.
More complex parameterizations
Beyond simple matrix or vector factorizations, a range of studies has investigated the optimization behavior of more complex architectures. For example, [16] analyze linear convolutional networks, while [1, 2] establish convergence guarantees for shallow and deep overparameterized networks, respectively. In the case of ReLU networks, [34] and others have characterized the gradient flow dynamics, highlighting convergence regimes. [25] provide convergence analyses for deep linear networks, showing that gradient-based optimization implicitly favors flatter minima. Similarly, [8] study global convergence of gradient descent in infinitely wide two-layer networks, showing how overparameterization can ensure convergence to near-optimal solutions.
@@ 1.2 Notation
For any positive integer $L$, the set of integers from $1$ to $L$ is denoted by $[L]$. The coordinate-wise product, also called the Hadamard product, of two vectors $x$ and $y$ in $\mathbb{R}^{p}$ is denoted $x\odot y$, where for each $i$ in $[p]$, $(x\odot y)_{i}=x_{i}y_{i}$. The vector $x^{\odot L}$ is defined as $\left(x^{L}_{i}\right)_{i\in[p]}$ and if $\left(x^{j}\right)_{j\in[L]}\in\left(\mathbb{R}^{p}\right)^{L}$, then $\bigodot_{j=1}^{L}x^{j}$ refers to $\left(\prod_{j=1}^{L}x^{j}_{i}\right)_{i\in[p]}$. We use $|\cdot|$ to denote the absolute value which can be defined for vectors in $\mathbb{R}^{p}$: for any $x\in\mathbb{R}^{p}$, $|x|=\left(|x_{i}|\right)_{i\in[p]}$. Similarly, the sign function denoted $\text{sign}(\cdot)$, the square root function $\sqrt{\cdot}$, the hyperbolic sine $\sinh$, the inverse hyperbolic sine arcsinh and the logarithm $\log$ are considered as functions from $\mathbb{R}^{p}$ to $\mathbb{R}^{p}$ that apply to each component. The Jacobian of $G=\left(G_{1},\ldots,G_{p}\right)$ with respect to $\theta\in\mathbb{R}^{k}$ is denoted by $J_{G}(\theta)$ and is defined as:
@@
The Hessian of $G$ is denoted by $\nabla^{2}G$. We denote by #\# the cardinality of a set, and by $\%$ the modulo operation. The vector $\mathbf{1}\in\mathbb{R}^{p}$ denotes the unitary vector equal to $\left(1\right)_{i\in[p]}$ and $\lfloor\cdot\rfloor$ represents the floor function. For a vector $v\in\mathbb{R}^{p}$, $\text{diag}(v)$ denotes the diagonal matrix in $\mathbb{R}^{p\times p}$ having the diagonal equal to $v$.
@@ 2 Deep Diagonal Linear Networks
In this section, we state several results on the trajectory of the variable $\theta$ when parameterized by a Deep Diagonal Linear Network and when its parameters are trained by Gradient Flow. More precisely, we overparameterize our system by letting
$\theta=\bigodot_{j=1}^{L}u^{j},$ (1)
where $L\geqslant 2$ is the number of layers of the network and each $u^{j}\in\mathbb{R}^{d}$ corresponds to the $j$-th layer, as illustrated in Figure 2. Note that for $L=2$, we recover the classical Diagonal Linear Network. We consider the training of the Deep Diagonal Linear Network by Gradient Flow, namely the minimization of
$\mathcal{L}\bigg(\bigodot_{j=1}^{L}u^{j}\bigg),$
which can be written as follows:
$\forall j\in[L],\penalty 10000\ \forall t\geqslant 0,\quad\frac{du^{j}(t)}{dt}+\nabla_{j}\mathcal{L}\left(\bigodot_{k=1}^{L}u^{k}(t)\right)=0,$ (2)
where $\left(\nabla_{j}\mathcal{L}\left(\bigodot_{k=1}^{L}u^{k}\right)\right)_{i}=\frac{\partial}{\partial u^{j}_{i}}\mathcal{L}\left(\bigodot_{k=1}^{L}u^{k}\right)$ for any $j\in[L]$ and $i\in[d]$. By simple computations, we get the following dynamical system:
$\forall j\in[L],\penalty 10000\ \forall t\geqslant 0,\quad\frac{du^{j}(t)}{dt}+\left(\bigodot_{k\neq j}u^{k}(t)\right)\odot\nabla\mathcal{L}(\theta(t))=0.$ (3)
@@
Next, we first derive several properties of the trajectories of the weights of the network, and then derive the main result on the dynamic of $\theta$. Convergence guarantees are stated in the last section, under the same assumption on the initialization.